How many gametes are possible for an individual that has the following genotype:

            AaBbCCddEeFF                                   (Please show your work)
Answer #1

Number of gametes = 2^n

Where, n is the number of heterozygous gene pairs

Here, 3 heterozygous gene pairs are present (Aa, Bb, Ee)

So, number of gametes = 2^3 = 8

Please rate.